A. Suppose a brick wall of a house is 4.00 inches thick and that the thermal conductivity of the brick is about 0.84 J/smK . The area is 240 ft2 and there is a 12F° difference between the inside and outside. Assume a heater pumps the same amount of energy lost via conduction in the opposite direction that the energy is being transferred via conduction, i.e. the heater pumps energy from the cold side to the hot side of the wall. If the heater is 30% efficient, how much energy does the heater actually use in one hour?

B. Suppose a brick wall of a house is 4.00 inches thick and that the thermal conductivity of the brick is about 0.84 J/smK . The area is 240 ft2 and there is a 12F° difference between the inside and outside.

How much does it cost to operate this heater for one hour under the conditions of the brick wall and temperature difference originally given? Typical energy cost is 15 cents per kW·hr.
Answer #1

(A) L=4 in = 0.1016 m K = 0.84 Jl smk A = 240ft? AT = 12° F = -11:11°C = Ihr = 3600 sec 2 = 22-2967m² Z H = - - КАДТ L H = -(.. H n X 100 E 30 = 7372974.44 X 100 E E = 24576581,5 J (B); E 24576581,5 J -7 E - 24576581-5 x 2-778 X10 E =6.83kw-hr fast 7